# Large-sample analysis of cost functionals for inference under the coalescent

### v1

@@ 2.2 The cost sequence and importance sampling
$p(n\mathbf{y}_{0}^{(n)})=\mathbb{E}_{p}\left[\mathbb{I}_{\{n\mathbf{Y}^{(n)}(0)=n\mathbf{y}_{0}^{(n)}\}}\right].$
@@
$p(n\mathbf{Y}^{(n)}(\tau^{(n)}))=\sum_{i=1}^{d}p(\mathbf{e}_{i})\mathbb{I}_{\{n\mathbf{Y}^{(n)}(K)=\mathbf{e}_{i}\}}$
is available explicitly and (2.2) corresponds to the importance weight from the importance sampling algorithm with proposal distribution $q$. Choosing a deterministic $k\leq n\|\mathbf{y}_{0}^{(n)}\|_{1}\leq\tau^{(n)}$ yields truncated algorithms, which will be useful for the asymptotic analysis of importance weights. They do not correspond to exact algorithms in practice because the factor $p(n\mathbf{Y}^{(n)}(k))$ is intractable, though further approximations have been used to enact bias-variance trade-off (Jasra et al., 2011).
@@
We interpret the ratio
$\frac{p(n\mathbf{y}\mid n\mathbf{y}-\mathbf{v})}{q(n\mathbf{y}-\mathbf{v}\mid n\mathbf{y})}$ (2.5)
as the one-step cost of choosing the proposal $q$ in place of the true distribution $p$ in the backward step from $\mathbf{y}\in\frac{1}{n}\mathbb{N}^{d}\setminus\{\bm{0}\}$ to $\mathbf{y}-\frac{1}{n}\mathbf{v}$, for each possible step $\mathbf{v}=\mathbf{e}_{j},\mathbf{e}_{j}-\mathbf{e}_{i},i,j=1,\dots,d$. Then, the importance sampling weights can be interpreted in terms of the cumulative cost of all the steps. More generally, we define the following cost-counting sequence.
@@
Let the positive function $c^{(n)}(\mathbf{v}\mid\mathbf{y})$, represent the one-step cost of a backward jump from $\mathbf{y}$ to $\mathbf{y}-\frac{1}{n}\mathbf{v}$, for $\mathbf{y}\in\frac{1}{n}\mathbb{N}^{d}\setminus\{\bm{0}\},\mathbf{v}=\mathbf{e}_{j},\mathbf{e}_{j}-\mathbf{e}_{i},i,j=1,\dots,d$. The sequence of cost-counting processes is defined as $C^{(n)}=\{C^{(n)}(k)\}_{k\in\mathbb{N}}\subset\mathbb{R}_{+},n\in\mathbb{N}$, where $C^{(n)}(k)$ is the cumulative cost of performing the steps $\mathbf{Y}^{(n)}(0),\dots,$ $\mathbf{Y}^{(n)}(k)$, i.e.
@@
Note that the function $c^{(n)}$ can be of the form (2.5), for an arbitrary proposal $q$ whose support coincides with $p$, but it can also be more general. In the next section, we study first the cost $C^{(n)}$ in general. Then, in order to study the asymptotic behaviour of the normalised importance sampling weight $W^{(n)}$, the specific form (2.5) is used. The description of well-known specific proposals is postponed to Section 4, and the asymptotic analysis of the corresponding costs and weights is in Section 5.
@@ 3 Asymptotic analysis of the cost sequence
Consider the sequence $\mathbf{y}_{0}^{(n)}$ of samples of large size satisfying Assumption 1.1, and assume $\mathbf{Y}^{(n)}(0)=\mathbf{y}^{(n)}_{0}$. Furthermore, recall that naturally $M_{ij}^{(n)}(0)=0,\forall n\in\mathbb{N}$, $i,j=1,\dots d,$ and $C^{(n)}(0)=1,\forall n\in\mathbb{N}$.
In order to show convergence of the cost sequence of Definition 2.6, we will need the following assumption on the asymptotic behaviour of the cost of one step.
@@
There exist some continuous functions $a_{j},b_{ij},i,j=1,\dots,d,$ such that $b_{ij}\geq 1$, and
$\lim_{n\to\infty}\sup_{\mathbf{y}\in B_{\delta}^{(n)}}|n(c^{(n)}(\mathbf{e}_{j}\mid\mathbf{y})-1)-a_{j}(\mathbf{y})|=0\quad\text{and}\quad\lim_{n\to\infty}\sup_{\mathbf{y}\in B_{\delta}^{(n)}}|c^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}\mid\mathbf{y})-b_{ij}(\mathbf{y})|=0,$ (3.1)
for each $\delta>0$, where $B_{\delta}^{(n)}=\{\mathbf{y}\in\frac{1}{n}\mathbb{N}^{d}:y_{j}\geq\delta,j=1\dots,d\}$. This is equivalent to uniform convergence on compact sets in the state-space of the technical framework defined in Section 7.1.
Note that Assumption 3.2, which will be needed for the convergence of the cost sequence, requires knowledge of the first order approximation of the one-step cost of mutation steps and of the second order approximation of the one-step cost of coalescence steps.
@@
Let $\mathbf{Z}^{(n)}=(C^{(n)},\mathbf{Y}^{(n)},\mathbf{M}^{(n)})\subset\mathbb{R}_{+}\times\frac{1}{n}\mathbb{N}^{d}\setminus\{\bm{0}\}\times\mathbb{N}^{d^{2}},n\in\mathbb{N},$ be the sequence composed by the cost sequence $C^{(n)}$ of Definition 2.6, the scaled block-counting sequence $\mathbf{Y}^{(n)}$ of Definition 2.4 evolving backwards in time, and the mutation-counting sequence $\mathbf{M}^{(n)}$ of Definition 2.5, with initial conditions given by Assumption 1.1 and 3.1. Assume that the one-step costs satisfy Assumption 3.2. Fix $t\in[0,1)$. Then, as $n\to\infty,$ the sequence of processes $\tilde{\mathbf{Z}}^{(n)}=\{\mathbf{Z}^{(n)}(\lfloor{sn}\rfloor{})\}_{s\in[0,t]}$ converges weakly to the process $\mathbf{Z}=\{(C(s),\mathbf{Y}(s),\mathbf{M}(s))\}_{s\in[0,t]}\subset\mathbb{R}_{+}\times\mathbb{R}_{+}^{d}\times\mathbb{N}^{d^{2}}$, defined as follows. The state process $\mathbf{Y}=\{\mathbf{Y}(s)\}_{s\in[0,t]}$ is the deterministic process defined by
@@
[…]
@@ 3.1 Heuristic explanation of the convergence
In a single transition, the Markov chain $\mathbf{Z}^{(n)}$ goes from state $(c,\mathbf{y},\mathbf{m})\in\mathbb{R}_{+}\times\frac{1}{n}\mathbb{N}^{d}\setminus\{\bm{0}\}\times\mathbb{N}^{d^{2}}$ to state
  • $\left(c\,c^{(n)}(\mathbf{e}_{j}\mid\mathbf{y}),\mathbf{y}-\frac{1}{n}\mathbf{e}_{j},\mathbf{m}\right)$ with probability $\rho^{(n)}(\mathbf{e}_{j}|\mathbf{y})$;
  • $\left(c\,c^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}\mid\mathbf{y}),\mathbf{y}-\frac{1}{n}\mathbf{e}_{j}+\frac{1}{n}\mathbf{e}_{i},\mathbf{m}+\mathbf{e}_{ij}\right)$ with probability $\rho^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}|\mathbf{y})$,
where $\rho^{(n)}$ is described in Definition 2.2. This can be summarised in the following following operator $A^{(n)}$, which is the infinitesimal generator of $\tilde{\mathbf{Z}}^{(n)}$,
@@
$\quad+\sum_{i,j=1}^{d}\left[f\left(c\,c^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}|\mathbf{y}),\mathbf{y}-\frac{1}{n}\mathbf{e}_{j}+\frac{1}{n}\mathbf{e}_{i},\mathbf{m}+\mathbf{e}_{ij}\right)-f(c,\mathbf{y},\mathbf{m})\right]n\rho^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}|\mathbf{y}),$ (3.4)
@@
$\rho^{(n)}(\mathbf{e}_{j}|\mathbf{y}^{(n)})\xrightarrow[n\to\infty]{}\frac{y_{j}}{\|\mathbf{y}\|_{1}},\qquad n\rho^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}|\mathbf{y}^{(n)})\xrightarrow[n\to\infty]{}\lambda_{ij}(\mathbf{y}),\quad\quad i,j=1,\dots,d.$ (3.5)
@@
$\quad+\sum_{i,j=1}^{d}\left[f\left(c\,b_{ij}(\mathbf{y}),\mathbf{y},\mathbf{m}+\mathbf{e}_{ij}\right)-f(c,\mathbf{y},\mathbf{m})\right]\lambda_{ij}(\mathbf{y}).$ (3.6)
@@ 4.1 Griffiths–Tavaré (GT) proposal
Substituting the proposal $q_{\scriptscriptstyle GT}$ into (2.5) shows that the cost of a backward step from $\mathbf{y}\in\frac{1}{n}\mathbb{N}^{d}\setminus\{\bm{0}\}$ does not depend on the type of step, and, for $\mathbf{v}=\mathbf{e}_{j},\mathbf{e}_{j}-\mathbf{e}_{i},i,j=1,\dots,d$, it is equal to
@@
$c_{\scriptscriptstyle GT}^{(n)}(\mathbf{v}\mid\mathbf{y})=1-\frac{1}{n}\frac{d-1}{\|\mathbf{y}\|_{1}}+o\left(\frac{1}{n}\right),\quad\mathbf{v}=\mathbf{e}_{j},\mathbf{e}_{j}-\mathbf{e}_{i},i,j=1,\dots,d.$
@@ 4.2 Stephens–Donnelly (SD) proposal
$q_{\scriptscriptstyle SD}(\mathbf{n}-\mathbf{v}|\mathbf{n})=$$\begin{cases}\frac{n_{j}(n_{j}-1)}{\|\mathbf{n}\|_{1}(\|\mathbf{n}\|_{1}-1+\theta)}\frac{1}{\hat{\pi}[j|\mathbf{n}-\mathbf{e}_{j}]},&\text{ if }\mathbf{v}=\mathbf{\mathbf{e}}_{j},\quad j=1\dots d,\\ \frac{\theta P_{ij}n_{j}}{\|\mathbf{n}\|_{1}(\|\mathbf{n}\|_{1}-1+\theta)}\frac{\hat{\pi}[i|\mathbf{n}-\mathbf{e}_{j}]}{\hat{\pi}[j|\mathbf{n}-\mathbf{e}_{j}]},&\text{ if }\mathbf{v}=\mathbf{\mathbf{e}}_{j}-\mathbf{\mathbf{e}}_{i},\quad i,j=1\dots d,\\ 0,&\text{ otherwise},\end{cases}$ (4.2)
@@
$\hat{\pi}[j\mid\mathbf{n}]=\sum_{i=1}^{d}\frac{n_{i}}{\|\mathbf{n}\|_{1}+\theta}\sum_{m=0}^{\infty}\left(\frac{\theta}{\|\mathbf{n}\|_{1}+\theta}\right)^{m}(P^{m})_{ij},\quad j=1,\dots,d,$
@@
$c_{\scriptscriptstyle SD}^{(n)}(\mathbf{v}\mid\mathbf{y})=\frac{p(n\mathbf{y}\mid n\mathbf{y}-\mathbf{v})}{q_{\scriptscriptstyle SD}(n\mathbf{y}-\mathbf{v}\mid n\mathbf{y})}=\begin{cases}\hat{\pi}[j|n\mathbf{y}-\mathbf{e}_{j}]\frac{\|\mathbf{y}\|_{1}}{y_{j}},&\text{ if }\mathbf{v}=\mathbf{\mathbf{e}}_{j},\quad j=1\dots d,\\ \frac{\hat{\pi}[j|\mathbf{n}-\mathbf{e}_{j}]}{\hat{\pi}[i|\mathbf{n}-\mathbf{e}_{j}]}\frac{ny_{i}-1+\delta_{ij}}{ny_{j}},&\text{ if }\mathbf{v}=\mathbf{\mathbf{e}}_{j}-\mathbf{\mathbf{e}}_{i},\quad i,j=1\dots d,\\ 0,&\text{ otherwise}.\end{cases}$
@@
$\hat{\pi}[i\mid n\mathbf{y}-\mathbf{e}_{j}]=\frac{y_{i}}{\|\mathbf{y}\|_{1}}+\frac{1}{n}\frac{1}{\|\mathbf{y}\|_{1}}\left[\frac{y_{i}(1-\theta)}{\|\mathbf{y}\|_{1}}-\delta_{ij}+\sum_{i^{\prime}=1}^{d}\frac{y_{i^{\prime}}}{\|\mathbf{y}\|_{1}}\theta P_{i^{\prime}i}\right]+o\left(\frac{1}{n}\right),\quad i,j=1,\dots,d.$
@@
$c_{\scriptscriptstyle SD}^{(n)}(\mathbf{e}_{j}\mid\mathbf{y})=1+\frac{1}{n}\hat{a}_{j}(\mathbf{y})+o\left(\frac{1}{n}\right),\quad j=1,\dots,d,$
@@
$c_{\scriptscriptstyle SD}^{(n)}(\mathbf{e}_{i}-\mathbf{e}_{j}\mid\mathbf{y})=1+o(1),\quad i,j=1,\dots,d.$

### v2

@@ 2.2 The cost sequence and importance sampling
$p(n\mathbf{y}_{0}^{(n)})=\mathbb{E}_{p}\left[\mathbb{I}(n\mathbf{Y}^{(n)}(0)=n\mathbf{y}_{0}^{(n)})\right].$
@@
$p(n\mathbf{Y}^{(n)}(\tau^{(n)}))=\sum_{i=1}^{d}p(\mathbf{e}_{i})\mathbb{I}(n\mathbf{Y}^{(n)}(k)=\mathbf{e}_{i})$
is available explicitly and (2.2) corresponds to the importance weight from the importance sampling algorithm with proposal distribution $q$. Choosing $k<\tau^{(n)}$ yields truncated algorithms, which are inexact in practice because the factor $p(n\mathbf{Y}^{(n)}(k))$ is intractable. For our purposes, truncation is necessary for the asymptotic analysis of importance weights because our scaling limit will only describe algorithms while the number of remaining lineages is large. In Section 6 we show empirically that despite this limitation, our analysis yields insights about algorithms without truncation. In practice, truncated algorithms have been used for bias-variance trade-off by Jasra et al. (2011), who found that truncating at approx. 10% of the sample size yielded barely perceptible bias for simple mutation models and $n=100$. Griffiths and Tavaré (1994b) also investigated a variance reduction method in which importance sampling was carried out until 2 or 3 lineages remained, and the resulting sampling distribution was computed numerically. They found truncation at two remaining lineages to be optimal because of the computational cost of numerical methods at higher truncation levels.
@@
The product in the display above can be thought of as a cost as explained in the following. The product goes over the steps of the importance sampling algorithm, at each step the algorithm uses the proposal distribution $q$ in place of the true distribution $p$. In particular, we can interpret the ratio
$\frac{p(n\mathbf{y}\mid n\mathbf{y}-\mathbf{v})}{q(n\mathbf{y}-\mathbf{v}\mid n\mathbf{y})}=:c_{q}^{(n)}(\mathbf{v}\mid\mathbf{y})$ (2.5)
as a one-step cost of using the proposal $q$ in place of the true distribution $p$ in the backward step from $\mathbf{y}\in\frac{1}{n}\mathbb{N}^{d}\setminus\{\bm{0}\}$ to $\mathbf{y}-\frac{1}{n}\mathbf{v}$, for each possible step $\mathbf{v}\in\{\mathbf{e}_{j},\mathbf{e}_{j}-\mathbf{e}_{i}\},\;i,j\in\{1,\dots,d\}$. Then, the importance sampling weights can be interpreted in terms of the cumulative cost of all the steps. More broadly, for a general form of one-step cost, we define the following cost-counting sequence.
@@
For $n\in\mathbb{N}$, let $c^{(n)}(\mathbf{v}\mid\mathbf{y})$ be a positive function of $\mathbf{y}\in\frac{1}{n}\mathbb{N}^{d}\setminus\{\bm{0}\},\mathbf{v}\in\{\mathbf{e}_{j},\mathbf{e}_{j}-\mathbf{e}_{i}\},\;i,j\in\{1,\dots,d\}$, representing the one-step cost of a backward jump from $\mathbf{y}$ to $\mathbf{y}-\frac{1}{n}\mathbf{v}$. The sequence of cost-counting processes is defined as $C^{(n)}=\{C^{(n)}(k)\}_{k\in\mathbb{N}}\subset\mathbb{R}_{+},n\in\mathbb{N}$, where $C^{(n)}(k)$ is the cumulative cost of performing the steps $\mathbf{Y}^{(n)}(0),\dots,$ $\mathbf{Y}^{(n)}(k)$, i.e.
@@
In the next section, we study the cost $C^{(n)}$ for a general form of one-step cost functions $c^{(n)}$. Then, in order to study the asymptotic behaviour of the normalised importance sampling weight $W^{(n)}$ in (2.4), we choose $c^{(n)}$ to be of the form $c^{(n)}_{q}$, as in (2.5), for an arbitrary proposal $q$ whose support coincides with $p$. The description of well-known specific proposals is postponed to Section 4, and the asymptotic analysis of the corresponding costs and weights is in Section 5.
@@ 3 Asymptotic analysis of the cost sequence
Consider the sequence $n\mathbf{y}_{0}^{(n)}$ of samples of large size satisfying Assumption 1.1, and assume $\mathbf{Y}^{(n)}(0)=\mathbf{y}^{(n)}_{0}$.
Furthermore, recall that naturally we have $M_{ij}^{(n)}(0)=0,\forall n\in\mathbb{N}$, $i,j\in\{1,\dots d\},$ and $C^{(n)}(0)=1,\forall n\in\mathbb{N}$. In order to show convergence of the cost sequence of Definition 2.6, we will need the following assumption on the asymptotic behaviour of the cost of one step.
@@
There exist some continuous functions $a_{j},b_{ij},\;i,j\in\{1,\dots,d\},$ such that $b_{ij}\geq 1$, and
$\lim_{n\to\infty}\sup_{\mathbf{y}\in B_{\delta}^{(n)}}|n(c^{(n)}(\mathbf{e}_{j}\mid\mathbf{y})-1)-a_{j}(\mathbf{y})|=0\qquad\text{and}\qquad\lim_{n\to\infty}\sup_{\mathbf{y}\in B_{\delta}^{(n)}}|c^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}\mid\mathbf{y})-b_{ij}(\mathbf{y})|=0,$ (3.1)
for each $\delta>0$, where $B_{\delta}^{(n)}=\{\mathbf{y}\in\frac{1}{n}\mathbb{N}^{d}:y_{j}\geq\delta,j\in\{1,\dots,d\}\}$.
The convergence in Assumption 3.2 is equivalent to uniform convergence on compact sets in the state-space of the technical framework defined in Section 7.1. It will be needed for the convergence of the cost sequence, and requires knowledge of the first order approximation of the one-step cost of mutation steps as well as the second order approximation of the one-step cost of coalescence steps.
@@
Let $\mathbf{Z}^{(n)}=(C^{(n)},\mathbf{Y}^{(n)},\mathbf{M}^{(n)})\subset\mathbb{R}_{+}\times\frac{1}{n}\mathbb{N}^{d}\setminus\{\bm{0}\}\times\mathbb{N}^{d^{2}},n\in\mathbb{N},$ be the sequence composed by the cost sequence $C^{(n)}$ of Definition 2.6, the scaled block-counting sequence $\mathbf{Y}^{(n)}$ of Definition 2.4 evolving backwards in time, and the mutation-counting sequence $\mathbf{M}^{(n)}$ of Definition 2.5, with initial conditions given by Assumption 1.1 and 3.1. Assume that the one-step costs satisfy Assumption 3.2. Fix $t\in[0,1)$. Then, as $n\to\infty,$ the sequence of processes $\tilde{\mathbf{Z}}^{(n)}=\{\mathbf{Z}^{(n)}(\lfloor{sn}\rfloor{})\}_{s\in[0,t]}$ converges weakly to the process $\mathbf{Z}=\{(C(s),\mathbf{Y}(s),\mathbf{M}(s))\}_{s\in[0,t]}\subset\mathbb{R}_{+}\times\mathbb{R}_{+}^{d}\times\mathbb{N}^{d^{2}}$, defined as follows. The state process $\mathbf{Y}=\{\mathbf{Y}(s)\}_{s\in[0,t]}=\{(Y_{1}(s),\ldots,Y_{d}(s)\}_{s\in[0,t]}$ is the deterministic process defined by
@@
$C(s)$ $=\exp\left\{-\int_{0}^{s}\langle a(\mathbf{Y}(u)),d\mathbf{Y}(u)\rangle+\sum_{i,j=1}^{d}\int_{0}^{s}\log b_{ij}(\mathbf{Y}(u))dM_{ij}(u)\right\}$
$=\exp\left\{\sum_{i=1}^{d}y_{0,i}\int_{0}^{s}a_{i}\left(\mathbf{y}_{0}(1-u)\right)du\right\}\prod_{i,j=1}^{d}\prod_{k=1}^{M_{ij}(s)}b_{ij}\left(\mathbf{y}_{0}\left(1-T_{ij}^{k}\right)\right),$ (3.3)
@@ 3.1 Heuristic explanation of the convergence
The one-step transition probabilities of the Markov chain $\mathbf{Z}^{(n)}$, started from a state $(c,\mathbf{y},\mathbf{m})\in\mathbb{R}_{+}\times\frac{1}{n}\mathbb{N}^{d}\setminus\{\bm{0}\}\times\mathbb{N}^{d^{2}}$, are
$\mathbb{P}(\mathbf{Z}_{k+1}^{(n)}=(c^{\prime},\mathbf{y}^{\prime},\mathbf{m}^{\prime})|\mathbf{Z}_{k}^{(n)}=(c,\mathbf{y},\mathbf{m}))$
$=\begin{cases}\rho^{(n)}(\mathbf{e}_{j}|\mathbf{y})&\text{if }(c^{\prime},\mathbf{y}^{\prime},\mathbf{m}^{\prime})=(c\,c^{(n)}(\mathbf{e}_{j}\mid\mathbf{y}),\mathbf{y}-\frac{1}{n}\mathbf{e}_{j},\mathbf{m}),\\ \rho^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}|\mathbf{y})&\text{if }(c^{\prime},\mathbf{y}^{\prime},\mathbf{m}^{\prime})=(c\,c^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}\mid\mathbf{y}),\mathbf{y}-\frac{1}{n}\mathbf{e}_{j}+\frac{1}{n}\mathbf{e}_{i},\mathbf{m}+\mathbf{e}_{ij}),\end{cases}$
where $\rho^{(n)}$ is described in Definition 2.2. This can be summarised in the following operator $A^{(n)}$, which is the infinitesimal generator of $\tilde{\mathbf{Z}}^{(n)}$,
@@
$\qquad+\sum_{i,j=1}^{d}\left[f\left(c\,c^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}|\mathbf{y}),\mathbf{y}-\frac{1}{n}\mathbf{e}_{j}+\frac{1}{n}\mathbf{e}_{i},\mathbf{m}+\mathbf{e}_{ij}\right)-f(c,\mathbf{y},\mathbf{m})\right]n\rho^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}|\mathbf{y}),$ (3.4)
@@
$\rho^{(n)}(\mathbf{e}_{j}|\mathbf{y}^{(n)})\xrightarrow[n\to\infty]{}\frac{y_{j}}{\|\mathbf{y}\|_{1}},\hskip 18.49988ptn\rho^{(n)}(\mathbf{e}_{j}-\mathbf{e}_{i}|\mathbf{y}^{(n)})\xrightarrow[n\to\infty]{}\lambda_{ij}(\mathbf{y}),\qquad\qquad i,j\in\{1,\dots,d\}.$ (3.5)
@@
$\qquad+\sum_{i,j=1}^{d}\left[f\left(c\,b_{ij}(\mathbf{y}),\mathbf{y},\mathbf{m}+\mathbf{e}_{ij}\right)-f(c,\mathbf{y},\mathbf{m})\right]\lambda_{ij}(\mathbf{y}).$ (3.6)
@@ 4.1 Griffiths–Tavaré (GT) proposal
Substituting the proposal $q_{\scriptscriptstyle GT}$ into (2.5) shows that the cost of a backward step from $\mathbf{y}\in\frac{1}{n}\mathbb{N}^{d}\setminus\{\bm{0}\}$ does not depend on the type of step, and, for $\mathbf{v}\in\{\mathbf{e}_{j},\mathbf{e}_{j}-\mathbf{e}_{i}\},\;i,j\in\{1,\dots,d\}$, it is equal to
@@
$c_{\scriptscriptstyle GT}^{(n)}(\mathbf{v}\mid\mathbf{y})=1-\frac{1}{n}\frac{d-1}{\|\mathbf{y}\|_{1}}+o\left(\frac{1}{n}\right),\qquad\mathbf{v}\in\{\mathbf{e}_{j},\mathbf{e}_{j}-\mathbf{e}_{i}\},\;i,j\in\{1,\dots,d\}.$
@@ 4.2 Stephens–Donnelly (SD) proposal
$q_{\scriptscriptstyle SD}(\mathbf{n}-\mathbf{v}|\mathbf{n})=$ $\begin{cases}\frac{n_{j}(n_{j}-1)}{\|\mathbf{n}\|_{1}(\|\mathbf{n}\|_{1}-1+\theta)}\frac{1}{\hat{\pi}[j|\mathbf{n}-\mathbf{e}_{j}]},&\text{ if }\mathbf{v}=\mathbf{\mathbf{e}}_{j},\qquad j\in\{1\dots d\},\\ \frac{\theta P_{ij}n_{j}}{\|\mathbf{n}\|_{1}(\|\mathbf{n}\|_{1}-1+\theta)}\frac{\hat{\pi}[i|\mathbf{n}-\mathbf{e}_{j}]}{\hat{\pi}[j|\mathbf{n}-\mathbf{e}_{j}]},&\text{ if }\mathbf{v}=\mathbf{\mathbf{e}}_{j}-\mathbf{\mathbf{e}}_{i},\qquad i,j=\{1,\dots.d\},\\ 0,&\text{ otherwise},\end{cases}$ (4.2)
@@
$\hat{\pi}[j\mid\mathbf{n}]=\sum_{i=1}^{d}\frac{n_{i}}{\|\mathbf{n}\|_{1}+\theta}\sum_{m=0}^{\infty}\left(\frac{\theta}{\|\mathbf{n}\|_{1}+\theta}\right)^{m}(P^{m})_{ij},\qquad j\in\{1,\dots,d\},$
@@
$c_{\scriptscriptstyle SD}^{(n)}(\mathbf{v}\mid\mathbf{y})=\frac{p(n\mathbf{y}\mid n\mathbf{y}-\mathbf{v})}{q_{\scriptscriptstyle SD}(n\mathbf{y}-\mathbf{v}\mid n\mathbf{y})}=\begin{cases}\hat{\pi}[j|n\mathbf{y}-\mathbf{e}_{j}]\frac{\|\mathbf{y}\|_{1}}{y_{j}},&\text{ if }\mathbf{v}=\mathbf{\mathbf{e}}_{j},\qquad j\in\{1\dots d\},\\ \frac{\hat{\pi}[j|\mathbf{n}-\mathbf{e}_{j}]}{\hat{\pi}[i|\mathbf{n}-\mathbf{e}_{j}]}\frac{ny_{i}-1+\delta_{ij}}{ny_{j}},&\text{ if }\mathbf{v}=\mathbf{\mathbf{e}}_{j}-\mathbf{\mathbf{e}}_{i},\qquad i,j\in\{1,\dots,d\},\\ 0,&\text{ otherwise}.\end{cases}$
@@
$\hat{\pi}[i\mid n\mathbf{y}-\mathbf{e}_{j}]=\frac{y_{i}}{\|\mathbf{y}\|_{1}}+\frac{1}{n}\frac{1}{\|\mathbf{y}\|_{1}}\left[\frac{y_{i}(1-\theta)}{\|\mathbf{y}\|_{1}}-\delta_{ij}+\sum_{i^{\prime}=1}^{d}\frac{y_{i^{\prime}}}{\|\mathbf{y}\|_{1}}\theta P_{i^{\prime}i}\right]+o\left(\frac{1}{n}\right),\qquad i,j\in\{1,\dots,d\}.$
@@
$c_{\scriptscriptstyle SD}^{(n)}(\mathbf{e}_{j}\mid\mathbf{y})=1+\frac{1}{n}\hat{a}_{j}(\mathbf{y})+o\left(\frac{1}{n}\right),\qquad j\in\{1,\dots,d\},$
@@
$c_{\scriptscriptstyle SD}^{(n)}(\mathbf{e}_{i}-\mathbf{e}_{j}\mid\mathbf{y})=1+o(1),\qquad i,j\in\{1,\dots,d\}.$